# On the free metabelian lie-admissible algebras

### v1

@@ 1. Introduction
In recent years, algebras with metabelian identity have become popular objects in ring theory. Metabelian identity also can be stated as the solvability of index $2$. Various types of classical algebras with metabelian identity, such as Lie, Leibniz, Malcev, Jordan, etc. are considered. The variety of metabelian Lie algebras has attracted significant attention, see [16, 17, 18, 24]. A basis of the free metabelian Lie algebra was constructed in [1]. In an analogical way, a basis of the free metabelian Leibniz algebra was constructed in [7]. Symmetric polynomials in the free metabelian Lie algebras were considered in [8]. The generators of symmetric polynomials in free metabelian Leibniz algebras were found in [23]. A basis of the free metabelian Malcev algebra is constructed in [22]. For this reason, we add metabelian identity to a well-known class of algebras which are Novikov and Lie-admissible.
@@
It is well-known that, given a commutative algebra $A$ with a derivation $d$, the space $A$ under the product $x_{1}\circ x_{2}=x_{1}d(x_{2})$ forms a Novikov algebra. Moreover, every Novikov algebra can be embedded into an appropriate commutative algebra with derivation $d$ [4]. Using rooted trees, the monomial basis of the free Novikov algebra in terms of $\circ$ was constructed in [10]. In terms of Young diagrams, the basis was developed in [9]. By utilizing commutative algebra with derivation $d$ and a well-defined order, an alternative monomial basis of the free Novikov algebra is presented in [14]. The issues of solvability and nilpotency of Novikov algebras were addressed in [26]. In section $2$, we construct a basis of the free solvable Novikov algebra with an index of $2$. In section $3$, we explicitly describe the generators of the symmetric polynomials space in the free metabelian Novikov algebra.
Let’s shift our focus to metabelian Lie-admissible algebras. In the realm of Lie-admissible algebra theory, additional conditions like flexibility or power-associativity are crucial, leading to numerous noteworthy outcomes in this context, see [3, 20, 21]. Another important direction of the research on Lie-admissible algebras concerns the property that their associator satisfies relations defined by a natural action of the symmetric group of degree 3 [12, 13]. A basis of the free Lie-admissible algebra and the Gröbner-Shirshov base theory for Lie-admissible algebras is given in [5]. In addition, there is given an analogue of PWB-theorem for the pair of Lie and Lie-admissible algebras. A basis of the free Lie-admissible algebra in terms of commutator and anti-commutator is given in [19]. In section $4$, we construct the basis of a free metabelian Lie-admissible algebra in terms of commutator and anti-commutator. In Section $5$, we explicitly describe the symmetric polynomials for the multilinear part of a free metabelian Lie algebra.
@@ 2. Free metabelian Novikov algebra
An algebra is called metabelian Novikov if it satisfies the following identities:
@@
$(ab)(cd)=0.$ (3)
Let $X=\{x_{1},x_{2}\ldots\}$ be a set of generators. We denote by $\mathrm{Nov}\langle X\rangle$ and $\mathrm{M}\mathrm{Nov}\langle X\rangle$ the free Novikov and free metabelian Novikov algebra, respectively.
@@
where $n\geq 5$. For degree $3$, $\mathcal{N}_{3}$ is defined as follows:
@@
An algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$ satisfies to the following identities:
@@
The set $\bigcup\mathcal{N}_{i}$ is a linear basis of algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$.
@@
The statement for degrees $3$ and $4$ can be checked by straightforward calculations. For $n\geq 5$, firstly, we show that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ can be written as a sum of monomials from the set $\bigcup_{i\geq 5}\mathcal{N}_{i}$. By Lemma (2.1), we obtain that every monomial except $(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}$ equal to 00. By (2) and Lemma (2.1), we have
@@
Now, we consider an algebra $A\langle X\rangle$ with a basis monomials $\bigcup\mathcal{N}_{i}$ and multiplication ***. Let us define a multiplication on monomials $\bigcup\mathcal{N}_{i}$ in $A\langle X\rangle$ as follows:
@@ 3. Symmetric polynomials of the free metabelian Novikov algebra
Let $p(x_{1},x_{2},\ldots,x_{n})$ be a polynomial of the free metabelian Novikov algebra. The polynomial $p(x_{1},x_{2},\ldots,x_{n})$ is called symmetric if it satisfies the following condition:
$p(x_{1},x_{2},\ldots,x_{n})\sigma=p(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})=p(x_{1},x_{2},\ldots,x_{n}),$
@@
$p_{n}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<\ldots<j_{n-1}}(\cdots((x_{i}x_{j_{n-1}})x_{j_{n-2}})\cdots)x_{j_{1}},$
where $n\geq 5$.
@@
The space of symmetric polynomials in the free metabelian Novikov algebra is generated by the set $\mathcal{P}$.
@@
For simplicity, we denote $d(x)$ and $d^{n}(x)$ by $x^{\prime}$ and $(d^{n-1}(x))^{\prime}$, respectively. To describe generators of the space of symmetric polynomials of $\mathrm{Nov}\langle X\rangle$ in degree $3$, we need to find generators of the space of the symmetric polynomials of the differential commutative algebra of weight $-1$ in degree $3$. This is the monomials of the form
$\sum_{i=1}^{n}\sum_{j_{1}<j_{2}}x_{j_{1}}^{\prime}x_{j_{2}}^{\prime}x_{i}\;\;\textrm{and}\;\;\sum_{i=1}^{n}\sum_{j_{1}<j_{2}}x_{i}^{\prime\prime}x_{j_{1}}x_{j_{2}},$
which correspond to $p_{3,1}$ and $p_{3,2}$.
@@
which gives that we rewrite monomial $a^{\prime}b^{\prime}d^{\prime}c$ to $a^{\prime}b^{\prime}c^{\prime}d$. Finally, the remained monomials of weight $-1$ are $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{4}}^{\prime\prime\prime}$ and $x_{j_{1}}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}}^{\prime}$, where $i_{1}\leq i_{2}\leq i_{3}$ and $j_{1}\leq j_{2}\leq j_{3}\leq j_{4}$. By Theorem 2.2, these monomials are linearly independent, and generators of symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(d)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ are
$\sum_{j_{1}<j_{2}<j_{3}<j_{4}}x_{j_{1}}^{\prime}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}},\;\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<j_{3}}x_{i}^{\prime\prime\prime}x_{j_{3}}x_{j_{2}}x_{j_{1}},$
which correspond to $p_{4,1}$ and $p_{4,2}$.
By Theorem 2.2, starting from degree $5$, all monomials of $\mathrm{Com}\langle X\rangle^{(d)}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ except $x_{i}^{(n-1)}x_{j_{n-1}}\ldots x_{j_{1}}$ are equal to 00, where $x_{i}^{(n-1)}=(x_{i}^{(n-2)})^{\prime}$. Hence, starting from degree $5$, generators of symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(d)}/$$\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ are
$\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<\ldots<j_{n-1}}x_{i}^{(n-1)}x_{j_{n-1}}x_{j_{n-2}}\cdots x_{j_{1}},$
@@ 4. Free metabelian Lie-admissible algebra
Let us consider the polarization of metabelian Lie admissible algebra, i.e., we consider an algebra with two operations which is defined on metabelian Lie admissible algebra as follows:
@@
which hold in free metabelian Lie admissible algebra.
Let us construct a basis of the free metabelian Lie admissible algebra in terms of binary trees with two types of vertices $\bullet$ and $\circ$. We consider only trees of the following form:
$A=$******
@@
***$x_{i_{1}}$$x_{i_{2}}$$x_{i_{n-1}}$$x_{i_{n}}$
We place on vertices of the tree $\bullet$ and $\circ$ in all possible ways. On the leaves of the tree, we place generators from the alphabet $X$. This tree in a unique way corresponds to the sequence (*$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}}).$
@@
then $A$ correspond to $(\bullet_{x_{1}},\circ_{x_{2}},x_{3})$. We define a set of sequences as follows:
@@
2) If the rightmost vertex is white then the rightmost generator is bigger than the previous one, i.e., for $(\ldots,\circ_{i_{n-1}},x_{i_{n}})$, we have $i_{n-1}<i_{n}$;
@@
A set of monomials $\mathcal{M}$ is a basis of the free metabelian Lie-admissible algebra.
@@
5. Symmetric polynomials of the free metabelian Lie admissible algebra
For $n$, let us define the sequence (*$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})$, where *** can be $\circ$ or $\bullet$. For each sequence, we define a space as follows:
(*$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}=\sum_{i_{1},\ldots,i_{n}}x_{i_{1}}*(x_{i_{2}}*(\ldots(x_{i_{n-1}}*x_{i_{n}})\ldots)).$
@@ 5. Symmetric polynomials of the free metabelian Lie admissible algebra
For each space (*$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$, $n\geq 3$, we define a polynomial $p_{(*,*,\ldots,*,n)}$ as follows:
@@
$\mathrm{M}\mathrm{Lie}\langle X\rangle=\oplus_{(*,*,\ldots,*)}(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$
as a vector space, where $n\geq 3$. For example, if $n=3$ then
$\mathrm{M}\mathrm{Lie}\langle X\rangle=(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\bullet)}\oplus(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}\oplus\\ (\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\circ,\bullet)}\oplus(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\circ,\circ)}.$
Moreover, each monomial of the space (*$(*_{x_{i_{1}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ under action $S_{n}$ belongs to the same space. It remains to note that $p_{(*,*,\ldots,*,n)}$ is a symmetric polynomial, and for each space (*$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},$ $x_{i_{n}})_{(*,*,\ldots,*)}$ there is one unique symmetric polynomial which is $p_{(*,*,\ldots,*,n)}$. ∎

### v2

# On the free metabelian Novikov and metabelian Lie-admissible algebras

@@ 1. Introduction
In recent years, algebras with metabelian identity have become popular objects in ring theory. Metabelian identity also can be stated as the solvability of index $2$. Various types of classical algebras with metabelian identity, such as Lie, Leibniz, Malcev, Jordan, etc. are considered. The variety of metabelian Lie algebras has attracted significant attention, see [17, 22]. A basis of the free metabelian Lie algebra was constructed in [1]. In an analogical way, a basis of the free metabelian Leibniz algebra was constructed in [8]. Symmetric polynomials in the free metabelian Lie algebras were considered in [9]. The generators of symmetric polynomials in free metabelian Leibniz algebras were found in [21]. Other examples of Lie-admissible algebras are assosymmetric algebras [6]. A basis of the free metabelian Malcev algebra is constructed in [20]. For this reason, we add metabelian identity to a well-known class of algebras which are Novikov and Lie-admissible.
@@
It is well-known that, given a commutative algebra $A$ with a derivation $D$, the space $A$ under the product $x_{1}\circ x_{2}=D(x_{1})x_{2}$ forms a (right) Novikov algebra. Moreover, every Novikov algebra can be embedded into an appropriate commutative algebra with derivation $D$ [4]. Using rooted trees, the monomial basis of the free Novikov algebra in terms of $\circ$ was constructed in [11]. In terms of Young diagrams, the basis was developed in [10]. By utilizing commutative algebra with derivation $D$ and a well-defined order, an alternative monomial basis of the free Novikov algebra is presented in [15]. The issues of solvability and nilpotency of Novikov algebras were addressed in [24]. In section $2$, we construct a basis of the free solvable Novikov algebra with an index of $2$. In section $3$, we explicitly describe the symmetric polynomials for the multilinear part of the free metabelian Novikov algebra.
Let’s shift our focus to metabelian Lie-admissible algebras. In the realm of Lie-admissible algebra theory, additional conditions like flexibility or power-associativity are crucial, leading to numerous noteworthy outcomes in this context, see [3, 19]. Another important direction of the research on Lie-admissible algebras concerns the property that their associator satisfies relations defined by a natural action of the symmetric group of degree 3 [13, 14]. A basis of the free Lie-admissible algebra and the Gröbner-Shirshov base theory for Lie-admissible algebras is given in [5]. In addition, there is given an analogue of PBW-theorem for the pair of Lie and Lie-admissible algebras. A basis of the free Lie-admissible algebra in terms of commutator and anti-commutator is given in [18]. In section $4$, we construct the basis of a free metabelian Lie-admissible algebra in terms of commutator and anti-commutator. In Section $5$, we explicitly describe the symmetric polynomials for the multilinear part of the free metabelian Lie-admissible algebra.
@@ 2. Free metabelian Novikov algebra
An algebra is called metabelian (right) Novikov if it satisfies the following identities:
@@
$(ab)(cd)=0$ (3)
Let $X=\{x_{1},x_{2}\ldots\}$ be a countable set of generators. We denote by $\mathrm{Nov}\langle X\rangle$ and $\mathrm{M}\mathrm{Nov}\langle X\rangle$ the free Novikov and free metabelian Novikov algebra, respectively.
@@
where $n\geq 5$. We set
$\mathcal{N}_{1}=\{x_{i}\},\mathcal{N}_{2}=\{x_{i_{1}}x_{i_{2}}\}.$
For degree $3$, $\mathcal{N}_{3}$ is defined as follows:
@@
The algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$ satisfies the following identities:
@@
The set $\bigcup\mathcal{N}_{i}$ is a linear basis of the algebra $\mathrm{M}\mathrm{Nov}\langle X\rangle$.
@@
The statement for degrees less than 5 can be verified by direct calculations. For $n\geq 5$, firstly, we show that every monomial of $\mathrm{M}\mathrm{Nov}\langle X\rangle$ can be written as a sum of monomials from the set $\bigcup_{i\geq 5}\mathcal{N}_{i}$. By Lemma (2.1), we obtain that every monomial except $(\cdots((x_{i_{1}}x_{i_{2}})x_{i_{3}})\cdots)x_{i_{n}}$ is equal to 00. By (2) and Lemma (2.1), we have
@@
Now, we consider an algebra $A\langle X\rangle$ with a basis monomials $\bigcup\mathcal{N}_{i}$ and multiplication *. Let us define a multiplication on monomials $\bigcup\mathcal{N}_{i}$ in $A\langle X\rangle$ as follows:
@@ 3. Symmetric polynomials of the free metabelian Novikov algebra
Let $p(x_{1},x_{2},\ldots,x_{n})$ be a polynomial of the free metabelian Novikov algebra generated by a finite set $X=\{x_{1},x_{2},\ldots,x_{n}\}$. The polynomial $p(x_{1},x_{2},\ldots,x_{n})$ is called symmetric if it satisfies the following condition:
$\sigma p(x_{1},x_{2},\ldots,x_{n})=p(x_{\sigma(1)},x_{\sigma(2)},\ldots,x_{\sigma(n)})=p(x_{1},x_{2},\ldots,x_{n}),$
@@
$p_{n}=\sum_{i=1}^{n}\sum_{j_{1}<j_{2}<\ldots<j_{n-1}}(\cdots((x_{i}x_{j_{1}})x_{j_{2}})\cdots)x_{j_{n-1}},$
where $n\geq 5$. The multilinear part of the free metabelian Novikov algebra is a space consisting of all elements containing each $x_{i}$ exactly once.
@@
For the multilinear part of the free metabelian Novikov algebra, the symmetric polynomials have the form $\mathcal{P}$.
@@
For simplicity, we denote $D(x)$ and $D^{n}(x)$ by $x^{\prime}$ and $(D^{n-1}(x))^{\prime}$, respectively. For the multilinear part to describe the space of symmetric polynomials of $\mathrm{Nov}\langle X\rangle$ in degree $3$, we need to find the space of the symmetric polynomials of the differential commutative algebra of weight $-1$ in degree $3$. This is the monomials of the form
$x_{2}^{\prime}x_{1}^{\prime}x_{3}+x_{3}^{\prime}x_{2}^{\prime}x_{1}+x_{3}^{\prime}x_{1}^{\prime}x_{2}\;\textrm{and}\;x_{1}^{\prime\prime}x_{2}x_{3}+x_{2}^{\prime\prime}x_{1}x_{3}+x_{3}^{\prime\prime}x_{1}x_{2}.$
Rewriting the first polynomial by the operation of Novikov algebra, we obtain $p_{3,1}$. Rewriting the second polynomial, we obtain
$(x_{1}x_{2})x_{3}-x_{1}(x_{2}x_{3})+(x_{2}x_{1})x_{3}-x_{2}(x_{1}x_{3})+(x_{3}x_{1})x_{2}-x_{3}(x_{1}x_{2}).$
Adding to the last expression $p_{3,1}$, we obtain $p_{3,2}$.
@@
which gives that we rewrite monomial $a^{\prime}b^{\prime}d^{\prime}c$ to $a^{\prime}b^{\prime}c^{\prime}d$. Finally, the remained monomials of weight $-1$ are $x_{i_{1}}x_{i_{2}}x_{i_{3}}x_{i_{4}}^{\prime\prime\prime}$ and $x_{j_{1}}x_{j_{2}}^{\prime}x_{j_{3}}^{\prime}x_{j_{4}}^{\prime}$, where $i_{1}\leq i_{2}\leq i_{3}$ and $j_{1}\leq j_{2}\leq j_{3}\leq j_{4}$. By Theorem 2.2, these monomials are linearly independent, and for multilinear part symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}_{-1}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ are
$x_{1}^{\prime}x_{2}^{\prime}x_{3}^{\prime}x_{4}\;\textrm{and}\;x_{1}^{\prime\prime\prime}x_{2}x_{3}x_{4}+x_{2}^{\prime\prime\prime}x_{1}x_{3}x_{4}+x_{3}^{\prime\prime\prime}x_{1}x_{2}x_{4}+x_{4}^{\prime\prime\prime}x_{1}x_{2}x_{3},$
which correspond to $p_{4,1}$ and $p_{4,2}$, analogically as in degree $3$.
By Theorem 2.2, starting from degree $5$, all monomials of $\mathrm{Com}\langle X\rangle^{(D)}/\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ except $x_{i}^{(n-1)}x_{j_{n-1}}\ldots x_{j_{1}}$ are equal to 00, where $x_{i}^{(n-1)}=(x_{i}^{(n-2)})^{\prime}$. Hence, starting from degree $5$, symmetric polynomials of $\mathrm{Com}\langle X\rangle^{(D)}/$$\{a^{\prime\prime}b^{\prime}cd+a^{\prime}b^{\prime}c^{\prime}d\}$ of weight $-1$ are
$x_{1}^{(n-1)}x_{2}x_{3}x_{4}\cdots x_{n}+x_{2}^{(n-1)}x_{1}x_{3}x_{4}\cdots x_{n}+\\ x_{3}^{(n-1)}x_{1}x_{2}x_{4}\cdots x_{n}+\ldots+x_{n}^{(n-1)}x_{1}x_{2}x_{3}\cdots x_{n-1},$
@@ 4. Free metabelian Lie-admissible algebra
Let us consider the polarization of metabelian Lie-admissible algebra, i.e., we consider an algebra with two operations which is defined on metabelian Lie-admissible algebra as follows:
@@
which hold in free metabelian Lie-admissible algebra.
Let us construct a basis of the free metabelian Lie-admissible algebra in terms of binary trees with two types of vertices $\bullet$ and $\circ$. We consider only trees of the following form:
$A=$**
@@
*$x_{i_{1}}$$x_{i_{2}}$$x_{i_{n-1}}$$x_{i_{n}}$
We place on vertices of the tree $\bullet$ and $\circ$ in all possible ways. On the leaves of the tree, we place generators from the countable set $X$. This tree in a unique way corresponds to the sequence $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}}).$
@@
then $A$ corresponds to $(\bullet_{x_{1}},\circ_{x_{2}},x_{3})$. We define a set of sequences as follows:
@@
2) If the rightmost vertex is white then the rightmost generator is less than the previous one, i.e., for $(\ldots,\circ_{i_{n-1}},x_{i_{n}})$, we have $i_{n-1}\leq i_{n}$;
@@
The set of monomials $\mathcal{M}$ is a basis of the free metabelian Lie-admissible algebra.
@@
If the monomials $X_{1}$ and $X_{2}$ do not involve $\circ$ then we rewrite the product $X_{1}\bullet X_{2}$ according to the multiplication table of free metabelian Lie algebras.
@@
5. Symmetric polynomials of the free metabelian Lie-admissible algebra
For $n$, let us define the sequence $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})$, where * can be $\circ$ or $\bullet$. For each sequence, we define a space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ as follows:
$(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}=\sum_{i_{1},\ldots,i_{n}}x_{i_{1}}*(x_{i_{2}}*(\ldots(x_{i_{n-1}}*x_{i_{n}})\ldots)).$
@@ 5. Symmetric polynomials of the free metabelian Lie-admissible algebra
For each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$, $n\geq 3$, we define a polynomial $p_{(*,*,\ldots,*,n)}$ as follows:
@@
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle=\oplus_{(*,*,\ldots,*)}(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$
as a vector space, where $\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{\geq 3}\langle X\rangle$ is a multilinear part of the free metabelian Lie-admissible algebra of degree greater than $2$. For example, if $n=3$ then
$\mathrm{M}\mathrm{Lie}\textrm{-}\mathrm{adm}_{3}\langle X\rangle=(\bullet_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\bullet)}\oplus(\bullet_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\bullet,\circ)}\oplus\\ (\circ_{x_{i_{1}}},\bullet_{x_{i_{2}}},x_{i_{3}})_{(\circ,\bullet)}\oplus(\circ_{x_{i_{1}}},\circ_{x_{i_{2}}},x_{i_{3}})_{(\circ,\circ)}.$
Moreover, each monomial of the space $(*_{x_{i_{1}}},\ldots,*_{x_{i_{n-1}}},x_{i_{n}})_{(*,*,\ldots,*)}$ under action $S_{n}$ belongs to the same space. It remains to note that $p_{(*,*,\ldots,*,n)}$ is a symmetric polynomial, and for each space $(*_{x_{i_{1}}},*_{x_{i_{2}}},\ldots,*_{x_{i_{n-1}}},$ $x_{i_{n}})_{(*,*,\ldots,*)}$ there is one unique symmetric polynomial which is $p_{(*,*,\ldots,*,n)}$. ∎